OpenStudy (anonymous):

William has batted 150 times and has a 0.220 batting average. If William can manage a hit everytime he bats for the rest of the season, what is the LEAST number of times he must bat to raise his batting average to 0.300? (Note: batting average = # of hits/# of times batted)

OpenStudy (ghazi):

average hit= number of hits/ number of times batted =0.220... number of hits= 0.22*150=33 to make average of 0.3 number of time he should bat is 0.3=33/number of times batted number of times batted =33/0.3=110

OpenStudy (anonymous):

choices r: 5 12 18 45 51

OpenStudy (anonymous):

srry. 4got about those

OpenStudy (ghazi):

well...again i would say...my answer is correct...because increasing average depends upon number of hits

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

:)

OpenStudy (ghazi):

:)